OpenStudy (anonymous):

Do they have instructions on this site that shows you how to create neat fractions and other neat math symbols

OpenStudy (anonymous):

if you come to chat i can show you some

OpenStudy (anonymous):

i can't show it here because it will just show up in math

OpenStudy (anonymous):

I am in chat

OpenStudy (anonymous):

\[ frac{2}{4} \]

OpenStudy (anonymous):

need \ before frac

OpenStudy (anonymous):

\[ \frac{2}{4} \]

OpenStudy (anonymous):

i wrote some others in chat as examples. you can see _ gives subscript

OpenStudy (anonymous):

\[e^{-x}\]

OpenStudy (anonymous):

\[e^{-x}\]\[\int_1^x \frac{1}{t}dt=\ln(x)\]

OpenStudy (anonymous):

you cannot use e^-x because only the minus sign will be a superscript. that is why you need {-x}

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

same thing with subscripts

OpenStudy (anonymous):

you can use \log_b but you need \lim_{h \rightarrow 0}

OpenStudy (anonymous):

Thank you! I really appreciate it.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]\]

OpenStudy (anonymous):

if you ever get stuck just google what you want web resources for latex are endless

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right. \]

OpenStudy (anonymous):

fancy one, i looked it up

OpenStudy (anonymous):

Yeah, I am used to $$ to start it of with and was not sure what starts and ended it here

OpenStudy (anonymous):

it turns out that $ and $$ are not actually latex commands, at least so i read. i use them, especially $

OpenStudy (anonymous):

Yeah, word press uses them

OpenStudy (anonymous):

but i think it might actually be a hack not sure

OpenStudy (anonymous):

Thanks a lot you help me a lot.

OpenStudy (anonymous):

\[ m = \frac{y_{2}-y_{1}} {y_{2}-y_{1}} \]

OpenStudy (anonymous):

\[ m = \frac{y_{2}-y_{1}} {x_{2}-x_{1}} \]